What is the concentration of HN(CH3)3+ needed in a solution that is 0.27 M trimethylamine, (CH3)3N, to maintain a pH of 9.59? Use the acid/base equilibrium constants found in the appendix of the lab manual.
Answer #1

To solve this problem we use Henderson equation.

pOH = pkb + log ([Conj. Acid] / [ weak base])

pOH = pkb + log ([HN(CH3)3+] / [trimethyl amine])

pkb of trimethyl amine (weak base) = 4.20

we know pH = 14 - pOH

(14-pH) = 4.20 + log ([HN(CH3)3+] / [ 0.27 ] )

[HN(CH3)3+ ]= 0.44 M